OpenStudy (anonymous):

Help with Trig question? Having to do with Angle sum and difference formulas. Find the exact value of sin[ sin^-1(5/13)-cos^-1 (-7/25)]. Any help would be appreciated!

OpenStudy (bibby):

sin(a-b)=sin(a)cos(b)-cos(a)sin(b) \(\sin[ \arcsin(5/13)-\arccos (-7/25)]=\\ sin(\arcsin(5/13))cos(\arccos (-7/25))-cos(\arcsin(5/13))sin(\arccos (-7/25))\) now the left hand cancels out pretty evenly \(\sin[sin^{-1}(\dfrac{5}{13})]=\dfrac{5}{13}\) same thing for the -7/25 brb pee break

OpenStudy (anonymous):

Ok, so the answer is just the value for sin ^-1

OpenStudy (bibby):

ok back so now we have \( \sin[ \arcsin(5/13)-\arccos (-7/25)]=\\ sin(\arcsin(5/13))cos(\arccos (-7/25))-cos(\arcsin(5/13))sin(\arccos (-7/25)) \\\dfrac{5}{13}*\dfrac{-7}{25}-\cos(\arcsin(5/13))\sin(\arccos (-7/25))\) \(\cos(\arcsin(5/13))=\dfrac{12}{13}\)

OpenStudy (bibby):

that leaves us with \(\\\dfrac{5}{13}*\dfrac{-7}{25}-\dfrac{12}{13}\sin(\arccos (-7/25))\) one last thingie to simplify

OpenStudy (bibby):

we can use another right triangle here. \(7^2+24^2=25^2\) if arccos of our angle is -7/25

OpenStudy (bibby):

\(\\\dfrac{5}{13}*\dfrac{-7}{25}-\dfrac{12}{13}*\dfrac{24}{25}=\dfrac{-35}{325}-\dfrac{288}{325}=\dfrac{-35-288}{325}=-\dfrac{323}{325} \)

OpenStudy (anonymous):

Ohhh, I started with the triangles first, and got lost. Thank you so much for walking me through that!

OpenStudy (bibby):

no problem, pls buy me lunch or dinner or linner as payment

OpenStudy (bibby):

p.s. that was a joke, now I'm off to run some errands and get food, feel free to tag me in any of your future questions using " @bibby "

OpenStudy (anonymous):

Hahaha, I would if I wasn't a poor college student now. :p thanks @bibby